As part of a​ study, 800 college football players are randomly chosen and their weights taken. The distribution of the weights is approximately normal. The average weight is 235 pounds and the standard deviation is 25 pounds. Approximately how many players had weights between 235 and 252​ pounds?
Answer #1

Solution :

Given that ,

mean = \mu = 235

standard deviation = \sigma = 25

P(235 < x < 252) = P[(235 - 235)/ 25) < (x - \mu) /\sigma  < (252 - 235) / 25) ]

= P(0 < z < 0.68)

= P(z < 0.68) - P(z < 0)

Using z table,

= 0.7517 - 0.5

= 0.2517

= 800 * 0.2517 = 201.36

= 201 players.